# Moner: Motion Correction in Undersampled Radial MRI with Unsupervised Neural Representation

### v1

@@ 4.3 Ablation Studies
Table 3: Quantitative results of MR images by our Moner ablating the motion model on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5.
@@
Forward Model MR Image Motion
@@
Table 4: Quantitative results of MR images and motion parameters by our Moner with different forward models on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5.
@@
Figure 7: Quantitative results of motion parameters by our Moner with different hash encodings on a sample (##\##5) of the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5.
@@
Table 5: Quantitative results of MR images by our Moner with different hash encodings on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5.
@@ 5 Conclusion and Limitation
While the proposed Moner demonstrates promising MoCo performance, it has several limitations. First, our Moner is currently designed for 2D MRI, while 3D MRI MoCo is more practical, as subjec’s movements occur in 3D space. However, we emphasize that extending it to 3D is feasible with advanced INR architectures, such as $K$-plane (Fridovich-Keil et al., 2023) and Hexplane (Cao & Johnson, 2023). Additionally, our Moner primarily focuses on radial MRI by leveraging the naive Fourier-slice theorem. Adapting it to handle more diverse MRI sampling patterns (e.g., Cartesian and spiral) also presents a potential direction for future work.
6 Reproducibility Statement
This work ensures high reproducibility. We conduct all experiments on two public MRI datasets (fastMRI (Knoll et al., 2020) and MoDL (Aggarwal et al., 2018)) and provide detailed descriptions of the data pre-processing pipeline in Sec. 4.1. Furthermore, we implement the compared methods and evaluation metrics based on the original papers or publicly available code, with additional details in the Appendix A.1 and A.2.
@@ Appendix A Appendix
A.1 Additional Details of Metrics
@@
A higher value of the proposed metric denotes greater MoCo accuracy. The intuition behind this metric is that if the transformation between the GT and reconstructed images is linear, the errors in the predicted MoCo parameters across different spokes are identical, i.e., $\Delta\vartheta_{1}=\Delta\vartheta_{2}=\cdots=\Delta\vartheta_{n}$ and $\Delta\bm{\tau}_{1}=\Delta\bm{\tau}_{2}=\cdots=\Delta\bm{\tau}_{n}$, resulting in a standard deviation of 0. Compared to the conventional $\ell_{1}$ metric, our new metric provides a more accurate assessment of MoCo accuracy.
@@
A.2 Additional Details of Baselines
@@
A.3 Additional Visual Results

### v2

@@ 4.3 Ablation Studies
Table 3: Quantitative results of MR images by our Moner ablating the motion model on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5. Results of t-test statistical tests are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and (not significant, $p$-value $\geq$ 0.05).
@@
Table 4: Quantitative results of MR images and motion parameters by our Moner with different forward models on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5. Results of t-test statistical tests are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and ($p$-value $\geq$ 0.05).
Forw. Model MR Image Motion
@@
Table 5: Quantitative results of MR images by our Moner with different hash encodings on the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5. Results of t-test statistical tests are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and ($p$-value $\geq$ 0.05).
@@ 5 Conclusion and Limitation
Figure 8: Motion trajectory by our Moner with different hash encodings on a sample (##\##5) of the fastMRI dataset for AF = 2$\times$ and MR = $\pm$5.
@@
While the proposed Moner demonstrates promising MoCo performance, it has several limitations. First, our Moner is currently designed for 2D MRI, while 3D MRI MoCo is more practical, as subject’s movements occur in 3D space. Secondly, the quasi-static motion model assumes rigid motion between acquisition frames, but it falls short in simulating certain complex motion scenarios. These include motion occurring within acquisition frames, non-rigid motion, and intricate spin-history effects. However, we emphasize that extending it to 3D is feasible with advanced INR architectures, such as $K$-plane (Fridovich-Keil et al., 2023) and Hexplane (Cao & Johnson, 2023). Also, for non-rigid motion, Moner can be extended by incorporating deformation field modeling (Reed et al., 2021). Additionally, our Moner primarily focuses on radial MRI by leveraging the naive Fourier-slice theorem. Adapting it to handle more diverse MRI sampling patterns (e.g., Cartesian and spiral) also presents a potential direction for future work.
@@ Appendix A Appendix
A.1 An Extension of Our Moner for 3D Radial MRI
Figure 9: Illustration of the relationship between stack-of-radial k-space data and 2D projections in the spatial domain. The 2D inverse Fourier transform (IFT) of a spoke array (orange) from a specific view corresponds to the respective 2D projections (purple).
Method
Our Moner is currently designed for 2D radial MRI. Here, we demonstrate its extension to 3D radial MRI. Using a 3D stack-of-radial sampling (Feng, 2022), the 2D inverse Fourier transform of a spoke array from a specific view corresponds to the respective 2D projections, as illustrated in Fig. 9. This allows Moner to be extended to 3D radial MRI by solving a parallel back-projection problem in 3D space. Moreover, we learn 6 motion parameter $(\vartheta,\varphi,\psi,\tau_{x},\tau_{y},\tau_{y})$ for each motion state to account for subject’s movements occurring in 3D space. Specifically, a rotation matrix ${\bm{A}}(\vartheta,\varphi,\psi)\in\mathbb{R}^{3\times 3}$ and a shift vector $\bm{\tau}\in\mathbb{R}^{3}$ are defined as below:
${\bm{A}}(\vartheta,\varphi,\psi)$ $=\begin{pmatrix}\cos\psi&-\sin\psi&0\\ \sin\psi&\cos\psi&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}\cos\varphi&0&\sin\varphi\\ 0&1&0\\ -\sin\varphi&0&\cos\varphi\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&\cos\vartheta&-\sin\vartheta\\ 0&\sin\vartheta&\cos\vartheta\end{pmatrix},$ (10)
$\bm{\tau}$ $=\begin{pmatrix}\tau_{x}\\ \tau_{y}\\ \tau_{z}\end{pmatrix}.$
By integrating the rotation matrix ${\bm{A}}(\vartheta,\varphi,\psi)$ and shift vector $\bm{\tau}$ into the spatial transformation (Eq. 5), our Moner can effectively model and correct the rigid motion in 3D space.
Results on Simulated Data
To test the effectiveness of our Moner in 3D radial MRI, we conduct a simulation study on a 3D brain MR image with dimensions of 240$\times$240$\times$240 acquired by a 3T Siemens MRI scanner. We use a 3D stack-of-radial sampling pattern with the golden-angle acquisition scheme. Detailed parameters are as follows: FOV = 283$\times$283$\times$283, image spacing = 1$\times$1$\times$1 mm3, and total spoke views = 720. The AFs are set 2$\times$ and 4$\times$, corresponding to 360 and 180 views, respectively. We also simulate two level of rigid motion $\beta=\{5,10\}$ with random rotations of $\left[-\beta,\beta\right]^{\circ}$ and shifts $\left[-\beta,\beta\right]$ mm along the X-axis, Y-axis and Z-axis. Note that here we consider 120 motion states, i.e., our model totally estimates 120$\times$6 = 720 motion parameters.
Fig. 10 presents the qualitative comparisons. NuIFFT struggles to achieve satisfactory MRI reconstructions, producing noticeable artifacts. In contrast, our method produces results that are visually close to the GTs, preserving both global structures and local details. Quantitative results in Table 6 further validate the significant improvements of our method over traditional NuIFFT. Additionally, Fig. 10 shows that our Moner accurately estimates 3D motion parameters.
Figure 10: Qualitative results of 3D MR images (Axial, Sagittal, and Coronal views) by NuIFFT and our Moner on the 3D brain data for AF = 2×2$\times$2 ×, MR = $\pm$ 10 and AF = 3×3$\times$3 ×, MR = $\pm$ 10.
Table 6: Quantitative results of 3D MR image by NuIFFT and our Moner the 3D brain data. Results of t-test statistical tests comparing our Moner to NuIFFT are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and (not significant, $p$-value $\geq$ 0.05).
AF MR NuIFFT Moner (Ours)
PSNR SSIM PSNR SSIM
2$\times$ $\pm$5 23.64∗∗ 0.499∗∗ 31.58 0.844
$\pm$10 21.77∗∗ 0.435∗∗ 31.64 0.843
4$\times$ $\pm$5 23.29∗∗ 0.439∗∗ 30.68 0.810
$\pm$10 21.81∗∗ 0.391∗∗ 30.70 0.813
Figure 11: Qualitative results of motion parameter estimation by our Moner on the 3D brain data for AF = 4×4$\times$4 × and MR = $\pm$ 10.
Figure 12: On a real-world T1w human brain 3D MRI volume with dimensions of 120$\times$120$\times$120 and a voxel size of 2$\times$2$\times$2 mm3, acquired using a UIH uMR 790 scanner, our Moner effectively reduces motion artifacts and recovers both global structures and detailed image features.
Figure 13: Visualizations of motion trajectory estimated by our Moner on a real-world T1w human bran 3D MRI volume with dimensions of 120$\times$120$\times$120 and a voxel size of 2$\times$2$\times$2 mm3, acquired using a UIH uMR 790 scanner. Scanning a single spoke requires 7.1 ms. The whole acquisition (6,5000 spokes in total) thus takes about 461 s. The subject is instructed to make abrupt movements (head shaking motion) three times during the acquisition. The estimated trajectory is generally consistent with the motion pattern.
Results on Real-World Data
To evaluate the effectiveness of our Moner on real-world MRI data, we collect two T1w brain 3D scans from a single subject using a 3.0 T United Imaging Healthcare (UIH) uMR 790 scanner. To obtain motion-free data and motion-corrupted data, the subject is instructed to remain still or make abrupt movements (head shaking motion) three times during acquisition. The used radial Spoiled gradient echo sequence parameters we used are as follows: TR = 7.1 ms; TE = 3 ms; flip angle = 17; matrix size = 120$\times$120$\times$120; FOV = 240$\times$240$\times$240 mm3; number of spokes = 65,000. Note that the data acquisition is approved by the institutional review board. Our Moner model does not access any prior motion knowledge. It instead assumes 650 motion stages and thus estimates unique i.e., 650$\times$6 = 3,900 motion parameters.
Fig. 12 shows the visualizations. Due to the 3D reconstruction task, only NuIFFT is included as an compared method. The result of NuIFFT suffers severely from motion artifacts, losing nearly a lots of image details. In contrast, our Moner’s reconstruction closely resembles the motion-free reference in global structures and successfully recovers tissue details, such as the lateral ventricles, as highlighted in the zoomed-in view. Moreover, we shows the visualizations of motion trajectory estimated by our Moner in Fig. 13. The estimated trajectory is generally consistent with the motion pattern. In summary, this study preliminarily validates the effectiveness of our method on real-world MRI data.
A.2 Additional Details of Metrics
@@
A lower value of the proposed metric denotes greater MoCo accuracy. The intuition behind this metric is that if the transformation between the GT and reconstructed images is linear, the errors in the predicted MoCo parameters across different spokes are identical, i.e., $\Delta\vartheta_{1}=\Delta\vartheta_{2}=\cdots=\Delta\vartheta_{n}$ and $\Delta\bm{\tau}_{1}=\Delta\bm{\tau}_{2}=\cdots=\Delta\bm{\tau}_{n}$, resulting in a standard deviation of 0. Compared to the conventional $\ell_{1}$ metric, our new metric provides a more accurate assessment of MoCo accuracy.
@@
A.3 Additional Details of Baselines
@@
Table 7: Quantitative results (Mean±$\pm$±STD in SSIM) of MR images by compared methods on the fastMRI and MoDL datasets. Results of t-test statistical tests comparing our Moner to baselines are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and (not significant, $p$-value $\geq$ 0.05). Here the “AF” and “MR” represent acceleration rate and motion range, respectively. The best and second performances are highlighted in bold and underline, respectively.
Dataset AF MR Analy. Optim. Sup. Unsup.
NuIFFT TV DRN-DCMB Score-MoCo Moner (Ours)
fastMRI 2$\times$ $\pm$2 0.537$\pm$0.057∗∗ 0.860$\pm$0.090 0.836$\pm$0.102∗∗ 0.866$\pm$0.045∗∗ 0.915$\pm$0.068
$\pm$5 0.364$\pm$0.065∗∗ 0.859$\pm$0.090 0.716$\pm$0.104∗∗ 0.849$\pm$0.038∗∗ 0.914$\pm$0.067
$\pm$10 0.274$\pm$0.074∗∗ 0.859$\pm$0.091 0.624$\pm$0.118∗∗ 0.851$\pm$0.042∗∗ 0.915$\pm$0.067
$\pm$15 0.251$\pm$0.074∗∗ 0.859$\pm$0.091 0.532$\pm$0.102∗∗ 0.857$\pm$0.041∗∗ 0.913$\pm$0.071
4$\times$ $\pm$2 0.460$\pm$0.042∗∗ 0.613$\pm$0.202∗∗ 0.812$\pm$0.096∗∗ 0.813$\pm$0.052∗∗ 0.888$\pm$0.077
$\pm$5 0.308$\pm$0.065∗∗ 0.645$\pm$0.203∗∗ 0.715$\pm$0.116∗∗ 0.800$\pm$0.060∗∗ 0.879$\pm$0.078
$\pm$10 0.221$\pm$0.067∗∗ 0.639$\pm$0.210∗∗ 0.570$\pm$0.124∗∗ 0.809$\pm$0.055∗∗ 0.882$\pm$0.081
$\pm$15 0.206$\pm$0.066∗∗ 0.610$\pm$0.208∗∗ 0.535$\pm$0.112∗∗ 0.804$\pm$0.055∗∗ 0.881$\pm$0.079
MoDL 2$\times$ $\pm$2 0.513$\pm$0.048∗∗ 0.888$\pm$0.015∗∗ 0.793$\pm$0.021∗∗ 0.766$\pm$0.156∗∗ 0.952$\pm$0.006
$\pm$5 0.336$\pm$0.041∗∗ 0.887$\pm$0.016∗∗ 0.613$\pm$0.019∗∗ 0.716$\pm$0.185∗∗ 0.952$\pm$0.006
$\pm$10 0.256$\pm$0.041∗∗ 0.887$\pm$0.016∗∗ 0.549$\pm$0.040∗∗ 0.718$\pm$0.180∗∗ 0.952$\pm$0.007
$\pm$15 0.246$\pm$0.041∗∗ 0.886$\pm$0.015∗∗ 0.356$\pm$0.029∗∗ 0.795$\pm$0.043∗∗ 0.949$\pm$0.010
4$\times$ $\pm$2 0.420$\pm$0.040∗∗ 0.861$\pm$0.063∗∗ 0.696$\pm$0.033∗∗ 0.552$\pm$0.158∗∗ 0.938$\pm$0.011
$\pm$5 0.257$\pm$0.036∗∗ 0.830$\pm$0.106∗∗ 0.578$\pm$0.031∗∗ 0.628$\pm$0.114∗∗ 0.933$\pm$0.011
$\pm$10 0.197$\pm$0.035∗∗ 0.835$\pm$0.100∗∗ 0.382$\pm$0.035∗∗ 0.643$\pm$0.182∗∗ 0.932$\pm$0.013
$\pm$15 0.185$\pm$0.037∗∗ 0.811$\pm$0.115∗∗ 0.371$\pm$0.043∗∗ 0.643$\pm$0.168∗∗ 0.933$\pm$0.007
A.4 Additional Visual Results